OpenStudy (anonymous):

Find the area of the equilateral triangle with sides 10 cm.. Ans is supposed to 43.3cm^2 (according to the answer sheet) but no workings are given :(

OpenStudy (anonymous):

ar of equilateral triangle=root3/4a^2

OpenStudy (anonymous):

First, divide in one of his heights the triangle into two triangles with a right angle (90, 60, 30 will be their angles). Then, you have 10 cm of hypothenuse, and 5 of base. You want to know the height. 10^2=height^2-5^2 75=height^2 sqrt(75)=height Then height*5/2 *2 (because you have 2 triangles) height*5=sqrt(75)*5=25*(sqrt(3))=43.3012~ cm2

OpenStudy (anonymous):

 x^2+25=100 x=5(3)^1/2 Area=1/2*10*5(3)^1/2 =25(3)^1/2/2 =43.301cm^2